Sharp Company manufactures a product for which the following standards have been set: Standard Quantity or Hours 3feet ?hours Standard Price or Rate 5 per foot ? per hour Standard Cost $15 Direct materials Direct labor During March, the company purchased direct materials at a cost of S57,250, all of which were used in the production of 3,280 units of product. In addition, 4,900 hours of direct labor time were worked on the product during the month. The cost of this labor time was $38,350. The following variances have been computed for the month: Materials quantity variance Labor spending variance Labor efficiency variance $3,300 U $1,010 F S 160 F

A. Compute the actual cost per foot for materials for March

Actual Cost per foot

Compute the price variance and the spending variance. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance))

Price Variance Unfavorable
Spending Variance Unfavorable

For direct labor:

a. Compute the standard direct labor rate per hour. (Do not round intermediate calculations. Round your final answer to 2 decimal places.)

Standard direct labor rate per hour


b. Compute the standard hours allowed for the month’s production.

Standard Hours    hours

c. Compute the standard hours allowed per unit of product. (Round your answer to 1 decimal place.)

Standard hours hours per unit
Answer #1
Materials Quantity variance
(AQ used - Std Qty allowed)*SR
(x - 3280*3)*5 =3300U
5x - 49200 = 3300U
x =(3300+49200)/5
10500
a) actual cost = 57,250/10500
5.45 per foot answer
(
b) Materials price variance
(Actual rate - standard rate)*AQ purchased
(5.45 - 5)*10,500
4750 U answer
Spending variance = 3300+4750
8050 U answer
2a) labor spending varian 1,010 F
labor efficiency varian 160 F
labor price variance 850 F
labor rate variance = ( Actual rate - standard rate)*Actual hours
(38,350- x*4900) = 850
x = (38350+850)/4900
8
Standard hours 8 per hour
b) Standard hours =
labor efficiency variance
(Actual hours - standard hours allowed)*Standard rate
(4900 -x ) *8 = 160
39200 - 8x = 160
x                       = (39200+160)/8
4920
Standard hours allowed 4,920 hours
c) Standard hours allowed per unit of product
4920/3280
1.5
Standard hours allowed per unit of product 1.5 hours per unit